OpenStudy (anonymous):

An investor invested a total of $2,200 in two mutual funds. One earned 9% profit, while the other earned 2% profit. If the investors total profit was $72, how much was invested in each mutual fund?

OpenStudy (anonymous):

To solve this you want to use a system of equations. Look for words to help make those equations: two mutual funds =2200 so x+y=2200 since we are dealing with interest we will rename x and y to p1 and p2 \[p _{1}+p_{2}=2200\] The simple interest equation is: principal Amount*interest*time=Interest gained \[p*i*t=I\] to keep it simple we guess time was 1 year then plug in the other numbers given: "One earned 9% profit, while the other earned 2% profit" and "total profit was $72" \[p _{1}*.09+p _{2}.02=72\] Then you just solve one for p1 or p2 and plug it into the other to simplify out one answer. Plug in that answer to get the other answer and you are done.

OpenStudy (anonymous):

I got p1=7 and p2=65, is that correct?

OpenStudy (anonymous):

If so which amount was invested in 9% and which amount was invested in 2%?

OpenStudy (anonymous):

sorry it has taken so long to reply always try to plug the numbers back in to double check. in this case ill help and just say those are wrong. so if you are still here try working this out step by step here and ill let you know how your doing each step.

OpenStudy (anonymous):

R u still here to help?